OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)?

OpenStudy (anonymous):

I need to see the slope

OpenStudy (anonymous):

I will give you medal

OpenStudy (anonymous):

The slope is 6/3

OpenStudy (anonymous):

1-(-5)=6 4-1=3 so the slope would be 6 over 3 6/3

OpenStudy (anonymous):

A. 2 B. -4/5 C. 5/4 D. -2

OpenStudy (anonymous):

ahh ok let me redo it

OpenStudy (anonymous):

Its A. 2 -5-1=-6 1-4=-3 -6 divided by -3 is 2

OpenStudy (cwrw238):

its difference of y coordinates divided by difference of x coordainates -5- 1 / 1-4 = -6/-3 = 2

OpenStudy (anonymous):

So witch answer is it?

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Yw